6. What are the WACC weights for debt and equity for an unlevered firm? (this is...

6.

What are the WACC weights for debt and equity for an unlevered firm? (this is more of a foundation question, thank you!)
Answer #1

Weighted Average Cost of Capital or WACC of a company is helpful in assessing return an investor could get on an investment in that particular company.

The formula for calculating WACC is given by:

where, Re = cost of Equity, Rd = Cost of Debt

tc = corporate tax-rate

E = Value of the Firm's Equity, D = Value of firm's Debt

V = D+E = Value of the firm

An Unlevered firm has Zero Debt i.e., D = 0

So, V = E+D = E+0 = E

For an unlevered firm WACC weights are: E/V = 1 and D/V = 0

Therefore the WACC of an Unlevered firm becomes

So, the WACC of an Unlevered firm is equal to its cost of Equity i.e., WACCU = Re